### v1

# Single-Shot Magnetization Reversal in Ferromagnetic Spin Valves Enabled by Heat Control

Abstract
We study laser-induced ultrafast magnetization reversal in ferromagnetic spin valve by comparing the effect of a direct laser excitation and an ultrashort hot-electron pulse. A wedged Cu layer is grown on top of the spin valve in order to tune the energy transmission to the magnetic stack, for both optical and hot-electron pulses. We demonstrate single-pulse magnetization reversal of the free layer by a hot-electron pulse. The influence of laser fluence, Cu thickness ($t_{\mathrm{Cu}}$), and pulse duration is investigated in detail. These results suggest that free layer heating plays a significant role in magnetization reversal. This work contributes to the understanding of ultrafast magnetization reversal due to nonlocal heat and spin transport occurring under strongly out-of-equilibrium conditions.
preprint: APS/123-QED
Introduction. Ultrafast stimuli, such as femtosecond laser pulses, have been shown to bring magnetic materials into an out-of-equilibrium state, resulting in ultrafast demagnetization [1]. Moreover, certain ferrimagnet systems, such as gadolinium-transition metal (Gd-TM)-based alloys and multilayers, as well as MnRuGa, undergo an ultrafast magnetization reversal upon single femtosecond laser pulse excitation [2, 3, 4, 5, 6, 7]. In Gd-TM systems, ultrafast magnetization reversal is mediated by angular momentum transfer between the two antiferromagnetically exchange-coupled Gd and TM sublattices.
In 2018, Iihama et al. demonstrated the magnetization reversal of a ferromagnetic layer (FM) in a GdFeCo/Cu/FM spin-valve structure [8]. An angular momentum transfer from the ferrimagnetic alloy (GdFeCo) to the FM layer, mediated by a spin current through the Cu spacer layer, was shown to be responsible for the FM layer reversal. From these measurements, the reversal process was shown to be compatible with a spin current originating from the demagnetization of Gd [8, 9].
In 2023, Igarashi et al. demonstrated single laser pulse-induced subpicosecond magnetization reversal in Gd-free [Co/Pt]/Cu/[Co/Pt] spin valves [10, 11, 12, 13], typically used for current-induced spin-transfer torque switching. The switching mechanism in ferromagnetic spin valves was found to differ from that of Gd-based ones. Indeed, as described in Refs. [13, 14], the injection of opposite-sign spin current into the free layer must be taken into account in order to explain the spin valve magnetization switching from parallel to antiparallel alignment. Although the physical mechanism is still unclear, there are two possible scenarios to explain this mechanism based on spin-current-mediated switching. The spin current can be either (i) due to the free layer demagnetization, or (ii) due to the reference layer demagnetization [13]. In the former case, the ultrafast demagnetization of the free layer would generate a spin current that propagates through the Cu spacer, experiences a reflection and spin-flip at the interface with the reference layer, and finally returns to the free layer to switch its magnetization. While these experimental results can be qualitatively reproduced by numerical calculations considering this reflection mechanism [15], it is clear that such a simplistic model needs refinement.
One fundamental question, both for applications and for understanding the physical mechanism behind these results, is whether light is necessary or if other heat-inducing stimuli can produce similar results. Indeed, ultrashort hot-electron pulses, generated by shining light on a non-magnetic metallic layer, have been shown to efficiently carry heat and induce ultrafast demagnetization of an adjacent FM layer [16, 17, 18], as well as induce full magnetization reversal in ferrimagnetic GdFeCo [19, 20].
In this Letter, we investigate magnetization reversal using both femtosecond light pulses and hot-electron pulses in [Co/Pt]3/Co/Cu(10)/[Co/Pt]2/Cu($t_{\mathrm{Cu}}$)/capping heterostructures. The effects of light and hot electrons can be tuned by varying the Cu thickness ($t_{\mathrm{Cu}}$), the nature of the capping, and the side of the sample illuminated by the laser pulse.
Figure 1: Light-induced switching in /Ta(5)/Pt(4)/[Co(0.82)/Pt(1)]3/Co(0.82)/Cu(10)/[Co(0.49)/Pt(1)]2/Cu($t_{\mathrm{Cu}}$)/Ta(5). (a) Schematic illustration of sample structure and the laser geometry used for indirect optical excitation. An LED light source with a center wavelength of 628 nm was used for taking MOKE images. (b) MOKE hysteresis loops obtained by measuring the Kerr rotation $\theta_{K}$ as a function of H applied perpendicular to the film plane for $t_{\mathrm{Cu}}$ = 2, 33, 67, and 100 nm. (c) MOKE images obtained before and after irradiation of a single laser pulse for $t_{\mathrm{Cu}}$ = 2 nm and $F_{\mathrm{p}}=9.3$ mJ/cm2, starting from parallel (P) and antiparallel (AP) alignment. (d) Same as (c), for $t_{\mathrm{Cu}}$ = 100 nm and $F_{\mathrm{p}}=64$ mJ/cm2 .
Sample stack and experimental methods. All samples were grown by sputtering. The basic stack structure is Ta(5)/Pt(4)/[Co(0.82)/Pt(1)]3/Co(0.82)/Cu(10)/[Co(0.49)/Pt(1)]2/Cu($t_{\mathrm{Cu}}$)/Ta(5) deposited on glass substrate (thicknesses are in nm). All samples have the same spin valve structure with a 10-nm-thick Cu spacer between two [Co/Pt] multilayers. We define the bottom (i.e., closer to the substrate) and top (i.e., farther from the substrate) [Co/Pt] multilayers as the reference and free layers, respectively. The reference layer has a higher Curie temperature than the free layer. A Cu layer was deposited on top of the spin valve to tune the optical absorption, as demonstrated in previous studies [16, 17]. The thickness of the Cu layer $t_{\mathrm{Cu}}$ was varied from 2 to 100 nm using a wedge deposition method [21]. Figure 1(a) shows a schematic illustration of the sample stack and typical experimental geometry. A linearly-polarized laser pulse was used for our experiments. The laser pulse was generated from a Ti:Sapphire femtosecond laser source with a wavelength of 800 nm and a repetition frequency of 5 kHz. The pulse duration was varied from 50 fs to 10 ps. A light-emitting diode (LED) source with a center wavelength of 628 nm was used to capture magneto-optical Kerr effect (MOKE) images of the samples [8]. We call it a direct optical excitation when the system was pumped and probed from the substrate side, and an indirect optical excitation when we pumped from the capping layer and probed from the substrate side. We always probed from the substrate side to measure the magnetic response, for all Cu thicknesses. Figure 1(b) shows hysteresis loops obtained via static MOKE measurements with an external magnetic field applied perpendicular to the film. We confirmed a perpendicular easy axis in both layers and four magnetic configurations. In this study, we focus on the P and AP states indicated in Fig. 1(b). The coercive field of both the free and the reference layers remained the same regardless of $t_{\mathrm{Cu}}$.
Single-shot experiment by shining a laser pulse on the capping layer. First, we investigate single-shot switching by shining a laser pulse on the capping layer, as illustrated in Fig. 1(a). Figure 1(c) shows MOKE images taken before and after laser excitation starting from P (red) and AP (blue) states for $t_{\mathrm{Cu}}$ = 2 nm. For such a thin Cu layer, the laser pulse directly reaches the spin valve, leading to direct laser-induced demagnetization. In this case, a clear magnetization reversal from P to AP state (P-to-AP switching) is observed as reported in previous works [13, 14]. Figure 1(d) shows the results of the single-shot experiment for $t_{\mathrm{Cu}}$ = 100 nm, wherein the laser pulse cannot reach the spin valve. Nevertheless, P-to-AP switching is still observed, indicating that hot electrons generated from the capping layer and flowing into the free layer are sufficient to trigger P-to-AP switching in the ferromagnetic spin valves. In contrast, AP-to-P switching was not observed in the studied samples. As previous studies have shown that reducing the thickness of the Cu spacer improves AP-to-P switching [13, 14], we henceforth focus solely on P-to-AP switching.
Figure 2(a) summarizes threshold fluences $F_{\mathrm{th}}$ for P-to-AP switching ($F_{\mathrm{P}}$) and for multidomain state ($F_{\mathrm{MD}}$). The value of $F_{\mathrm{th}}$ is determined for each state by fitting the domain area as a function of various laser pulse energies, as described in the Supplementary Material of Ref. [14]. Both $F_{\mathrm{P}}$ and $F_{\mathrm{MD}}$ are observed to increase with $t_{\mathrm{Cu}}$. The change in slope of $F_{\mathrm{P}}$ defines three different regimes (1, 2, and 3), as indicated in Fig. 2(a). The existence of these regimes can be associated to differences in energy absorption, which is calculated using a transfer matrix method and refractive indices for each material collected from Refs. [22, 23]. Optical energy absorption in the free layer, shown in Fig. 2(b), decreases significantly with increasing $t_{\mathrm{Cu}}$ and reaches less than 0.1% at $t_{\mathrm{Cu}}$ = 60 nm. From both experiments and calculations, we identify three distinct regions. In region 1 ($t_{\mathrm{Cu}}<30\,\mathrm{nm}$), direct optical excitation is the dominant mechanism. In region 2 ($30\,\mathrm{nm}<t_{\mathrm{Cu}}<70\,\mathrm{nm}$), both direct optical and indirect (hot electron) contributions are involved in the switching process. In region 3 ($t_{\mathrm{Cu}}>70\,\mathrm{nm}$), indirect excitation is the dominant mechanism. The slope of region 3 is observed to be similar to that of region 1, which contradicts the behavior predicted by the calculation in Fig. 2(b). One possible explanation is that the high phonon temperature induced by the laser pulse shortens the mean free path of electrons [18], resulting in fewer electrons reaching the spin valve. Consequently, greater laser power is required for magnetization reversal in thicker films.
We also investigate how the capping layer affects P-to-AP switching by testing different capping materials: Ta, Pt, and no capping. The lowest $F_{\mathrm{P}}$ is observed at approximately $t_{\mathrm{cap}}$ = 7 to 10 nm, which represents the optimal compromise between laser absorption depth and hot-electron scattering within the capping layer itself [17]. For further details on capping layer dependence, see End Matter.
Figure 2: Summary of the single-shot switching experiments in Ta(5)/Pt(4)/[Co(0.82)/Pt(1)]3/Co(0.82)/Cu(10)/[Co(0.49)/Pt(1)]2/Cu($t_{\mathrm{Cu}}$)/Ta(5). (a) Evolution of the threshold fluence for P-to-AP switching ($F_{\text{P}}$) and multidomain switching ($F_{\text{MD}}$) as a function of $t_{\mathrm{Cu}}$. (b) Calculation of laser energy absorption in the free and capping layers as a function of $t_{\mathrm{Cu}}$.
Figure 3: Pulse duration dependence of P-to-AP switching induced by indirect optical excitation. (a) MOKE images obtained after laser pulse irradiation for various pulse durations in Ta(5)/Pt(4)/[Co(0.82)/Pt(1)]3/Co(0.82)/Cu(10)/[Co(0.49)/Pt(1)]2/Cu(100)/Ta(5). We shine a laser pulse on the capping layer and a hot-electron pulse stimulates the switching. (b) Evolution of the threshold fluence for P-to-AP switching ($F_{\text{P}}$) and free-layer/reference-layer multidomain switching ($F_{\text{MD}}$) as a function of $\tau_{\mathrm{pulse}}$.
In the following, we investigate the effect of pulse duration $\tau_{\mathrm{pulse}}$ on P-to-AP switching threshold fluence, which can provide valuable insight into all-optical switching (AOS) [24, 21]. In a previous study, on a similar [Co/Pt]/Cu(10)/[Co/Pt] spin valve, P-to-AP switching was observed up to $\tau_{\mathrm{pulse}}$ = 1 ps, while a multidomain state in the free layer was observed instead for longer pulses [13]. The Cu layer on top of GdFeCo has been demonstrated to act as a heat sink, preventing long-term heat accumulation in GdFeCo and thus extending the pulse duration over which AOS can be observed [25]. From this, we expect P-to-AP switching to occur for extended pulse durations in our samples as well. Figure 3(a) shows MOKE images for various pulse durations taken after shining a laser pulse from the capping layer in Ta(5)/Pt(4)/[Co(0.82)/Pt(1)]3/Co(0.82)/Cu(10)/[Co(0.49)/Pt(1)]2/Cu(100)/Ta(5). P-to-AP switching can be observed with pulses up to $\tau_{\mathrm{pulse}}$ = 8 ps, due to the Cu heat sink. Furthermore, a multidomain state in the free layer appears in the center of the beam for longer pulse durations, such as 7, 8, and 10 ps. Figure 3(b) summarizes $F_{\mathrm{P}}$ and $F_{\mathrm{MD}}$, the latter for both free and reference layer. Both $F_{\mathrm{P}}$ and $F_{\mathrm{MD}}$ increase with pulse duration, as previously reported in Ref. [13] in samples that did not include a heat sink [13]. In our work, we confirm that incorporating a Cu heat sink results in an extended switching window to observe P-to-AP switching, while also demonstrating that the heat sink works in ferromagnetic spin valves similar to the previously reported GdFeCo/Cu system [25].
Figure 4: Single-shot experiment in Ta(5)/Pt(4)/[Co(0.82)/Pt(1)]3/Co(0.82)/Cu(10)/[Co(0.49)/Pt(1)]2/Cu($t_{\mathrm{Cu}}$)/Ta(5). (a) Schematic illustration of sample structures and the laser geometry used for direct excitation. We shine a laser pulse on the reference layer. MOKE images were obtained following irradiation of a single laser pulse for $t_{\text{Cu}}$ = 2 and 25 nm. The white region corresponds to the multidomain state in both layers. (b) Evolution of the threshold fluence for P-to-AP switching ($F_{\text{P}}$) and the multidomain state ($F_{\text{MD}}$) as a function of $t_{\text{Cu}}$.
Single-shot experiment shining on the reference layer. In order to compare the effect of direct and indirect excitation on P-to-AP switching in the same sample, we shine a laser pulse on the substrate side as illustrated in Fig. 4(a). Figure 4(b) shows MOKE images obtained after laser excitation for samples with $t_{\mathrm{Cu}}$ = 2 and 25 nm. P-to-AP switching is observed for $t_{\mathrm{Cu}}$ = 2 nm. Conversely, no P-to-AP switching is observed for $t_{\mathrm{Cu}}$ = 25 nm, despite having observed P-to-AP switching in the same sample with a laser excitation from the capping side. Figure 4(c) shows that $F_{\mathrm{P}}$ increases with $t_{\mathrm{Cu}}$ while $F_{\mathrm{MD}}$, for both free and reference layer, appears to stay constant. P-to-AP switching could be observed for $t_{\mathrm{Cu}}$ up to 20 nm.
To further characterize the switching mechanism, we inserted an insulating MgO layer between the free layer and 100-nm-thick Cu layer, as shown in Fig. 5(a). MgO suppresses electron transport, therefore blocking both spin and heat currents [26, 27, 28]. Figure 5(b) shows MOKE images taken after laser pulse irradiation in samples without and with the MgO layer. As presented earlier in Fig. 4(a) and shown again in Figure 5(b) (left), no P-to-AP switching is observed in this sample (it should be noted that the blue ring in Fig. 5(b) (top left) does not correspond to switching, as indicated by the line profile shown on the right side). Conversely, the same sample with an inserted MgO layer shows P-to-AP switching. Figure 5(c) summarizes $F_{\mathrm{th}}$ as a function of MgO thickness $t_{\mathrm{MgO}}$, where it can be observed that $F_{\mathrm{th}}$ does not depend on $t_{\mathrm{MgO}}$.
For further understanding, we perform a numerical calculation based on a two-temperature model [29, 18]. The calculation reveals that the electronic temperature of the free layer with the MgO layer is higher than that without MgO. In contrast, the electron temperature of the reference layer remains constant, both with and without the MgO layer. These results suggest an accumulation of electrons in the free layer due to the MgO layer, bringing heat and thus increasing the electronic temperature as illustrated in Fig. 5(a). This efficient rise in electronic temperature may facilitate the P-to-AP switching process. The details of the calculation are described in End Matter.
Figure 5: Single-shot experiment in Ta(5)/Pt(4)/[Co(0.82)/Pt(1)]3/Co(0.82)/Cu(10)/[Co(0.49)/Pt(1)]2/MgO($t_{\mathrm{MgO}}$)/Cu(100)/Ta(5). (a) Schematic illustration of sample structure and the laser geometry. We shine a laser pulse on the reference layer. (b) MOKE images obtained after irradiation of a single laser pulse and line profiles for $t_{\mathrm{MgO}}$ = 0 and 1 nm, starting from both P and AP states. (c) Evolution of threshold fluences for P-to-AP switching ($F_{\text{P}}$) and the multidomain state ($F_{\text{MD}}$) as a function of $t_{\text{MgO}}$. (d) MOKE images obtained after a laser pulse irradiation from the capping layer for $t_{\mathrm{MgO}}$ = 1 nm with $F_{\mathrm{p}}=137$ mJ/cm2.
We also investigate the effect of MgO insertion on switching stimulated by irradiation from the capping side. Figure 5(d) shows MOKE images after laser excitation for $t_{\mathrm{MgO}}$ = 1 nm. No P-to-AP switching is observed in spin valves with 1-nm-thick MgO insertion, even though a multidomain of the free layer is observed with high fluence $F_{\mathrm{p}}>100$ mJ/cm2, indicating that 1-nm-thick MgO strongly suppresses the electron transport. Free layer multidomain may be attributed to heat transport via phonon interaction in the MgO layer, resulting in the long-term demagnetization of the free layer [18, 30].
Here, we demonstrate that heat accumulation in the free layer plays a crucial role in P-to-AP switching. Nevertheless, the mechanism by which P-to-AP switching is observable in the absence of MgO with capping layer irradiation remains unclear. Given the thermal gradient present in this configuration, one possible explanation is that the free layer is excited more efficiently, resulting in a higher electron temperature than when irradiated from the substrate side.
Conclusion. We demonstrate single-shot magnetization switching in Co/Pt spin valves without direct laser excitation. Additionally, by incorporating a Cu heat sink layer, we can extend the pulse duration window for observing P-to-AP switching. Our comprehensive experiments reveal a critical requirement for the P-to-AP switching: the complete demagnetization of the free layer induced by ultrafast stimuli, such as laser pulses or hot-electron pulses. These findings provide deeper insight into ultrafast magnetization reversal driven by non-local spin transport under strongly out-of-equilibrium conditions.
Acknowledgements. We thank E. Díaz for fruitful discussion. This work is supported by the French National Research Agency (ANR) through the France 2030 government PEPR Electronic grants EMCOM (ANR-22-PEEL-0009) by the ANR SPOTZ project ANR- 20-CE24- 0003, SLAM ANR- 23-CE30-0047, the project MAT-PULSE from “Lorraine Université d’Excellence” reference ANR-15- IDEX-04-LUE, the Institute Carnot ICEEl, the Région Grand Est, the Metropole Grand Nancy, the “FEDERFSE Lorraine et Massif Vosges 2014-2020”, a European Union Program, the Academy of Finland (Grant No. 316857), the Sakura Program, the JSPS Bilateral Program, JSPS KAKENHI JP 24K22964, the Tohoku University-Universite de Lorraine Matching Funds, and CSIS cooperative research project in Tohoku University. This article is based upon work from COST Action CA23136 CHIROMAG, supported by COST (European Cooperation in Science and Technology). K. I. acknowledges Grant-in-Aid for JSPS Fellow (22J22178), X-NICS, and GP-Spin at Tohoku University. J.I. acknowledges support from JSPS Overseas Research Fellowships. All funding was shared equally among all authors.